OpenStudy (anonymous):

A ball is thrown vertically upward, what is the formula to find out how high it rises?

OpenStudy (anonymous):

let intial velocity be v then equating cahnges in potential nad kinetic energies we get 1/2mv^2 =mgh h=v^2/2g

OpenStudy (anonymous):

v=u-gt \[v ^{2}=u ^{2}+2as\] where v=0